U.S. U.K.
Wheat (bushels/ labor-hour) 4 2
Cloth (yards/ labor- hour) 2 1

a. express the cost of producing wheat and cloth in the U.S. and in the U.K. in terms of labor content

b. find Pc/Pw for both the U.S. and the U.K. in the absence of trade

c. express the limits for mutually advantageous trade in terms of Pc/Pw. What does this result indicate?
Answer #1

Solution:

a. Since US produces 4 bushels of wheat by using 1 labor-hour, cost of wheat in US in terms of labor = 1/4 labor-hour

Similarly, cost of wheat and cloth in terms of labor is calculated in the table:

US UK
Wheat (cost per bushel) 1/4 1/2
Cloth (cost per yard) 1/2 1

b. For US, PC/PW = (1/2)/(1/4) = 2

For UK, PC/PW = 1/(1/2) = 2

c. Although US has absolute advantage in the production of wheat and cloth but it has no comparative advantage. Similarly UK has absolute disadvantage in the production of wheat and cloth but it has no comparative advantage. Hence, no mutually advantageous trade is possible.